# Learning to combine top-down context and feed-forward representations under ambiguity with apical and basal dendrites

### v2

@@ 4.2 Contextual integration model architecture
The contextual integration model consists of three components: the pretrained backbone that computes representations of input images, the readout network for classifying their respective representations, and the topdown network which modulates the intermediate representations from the pretrained backbone based on contextual information. The backbone is implemented as a convolutional variational autoencoder (VAE), and was pretrained on unambiguous characters with the standard ELBO loss. VAEs are known to learn a smooth latent space, which we assume here to be an important ingredient to allow learning of top-down contextual modulations. Next, we freeze the VAE weights and train a readout (MLP) on the VAEs latent representations of images with a classification objective (cross entropy loss). Finally, we freeze these components and train the topdown network in two contextual tasks, with oracle and temporal context respectively, as described in the next sections. The topdown network $g(\cdot)$ is implemented as a multilayer perceptron (MLP) mapping the concatenated VAE latents and context vector to apical activations (eq. 3), which then modulate the VAEs intermediate representation. Various modulation functions (hadamard, additive, concatenation) are explored in the main text.
@@ 4.2.2 Temporal context
In the temporal case, the information is presented in a sequence, and preceding elements determine the contextual information available to the model at the current time $t$. Therefore, we leverage a Gated Recurrent Unit (GRU) network (Cho et al., 2014) to dynamically capture and represent the temporal dependencies in the sequence. We trained our GRU to predict the modulo 10 sum of MNIST digits in a sequence of two digits’, reflecting the task in our sequential MNIST/EMNIST dataset. An important detail is that the input to the GRU at time $t$ is not the image $x_{t}$ , but rather the latent $\mathbf{Ub}_{t}$, encoded by the pretrained backbone. After processing a sequence of digits ($x_{t-2},x_{t-1}$), the output state $o_{t}$ of the GRU was then used as a top-down context signal, in place of the one-hot context from the oracle case. We used a hidden state of size 128 for the GRU.
@@ 4.3.1 Loss function
The topdown network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function jointly across the following 5 conditions: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra). We use mean squared error as the loss criterion between the predicted (computed as in equation 2) and the target latent representations. Each minibatch is sampled from dataset $D$, and training data for each condition are equally balanced in the minibatch.
@@
3:for triplet $\texttt{triplet minibatch}(\mathbf{x_{0}},\mathbf{y_{0}}),\mathbf{x_{ambig}},(\mathbf{x_{1}},\mathbf{y_{1}})\texttt{ sampled from }D$ do

### v3

@@ 4.2 Contextual integration model architecture
The contextual integration model consists of three components: the pre-trained backbone that computes representations of input images, the readout network for classifying their respective representations, and the top-down network which modulates the intermediate representations from the pre-trained backbone based on contextual information. The backbone is implemented as a convolutional variational autoencoder (VAE), and was pre-trained on unambiguous characters with the standard ELBO loss. VAEs are known to learn a smooth latent space, which we assume here to be an important ingredient to allow learning of top-down contextual modulations. Next, we freeze the VAE weights and train a readout (MLP) on the VAEs latent representations of images with a classification objective (cross entropy loss). Finally, we freeze these components and train the top-down network in two contextual tasks, with oracle and temporal context respectively, as described in the next sections. The top-down network $g(\cdot)$ is implemented as a multi-layer perceptron (MLP) mapping the concatenated VAE latents and context vector to apical activations (eq. 3), which then modulate the VAEs intermediate representation. Various modulation functions (Hadamard, additive, concatenation) are explored in the main text.
@@ 4.2.2 Temporal context
In the temporal case, the information is presented in a sequence, and preceding elements determine the contextual information available to the model at the current time $t$. Therefore, we leverage a Gated Recurrent Unit (GRU) network (Cho et al., 2014) to dynamically capture and represent the temporal dependencies in the sequence. We trained our GRU to predict the modulo 10 sum of MNIST digits in a sequence of two digits’, reflecting the task in our sequential MNIST/EMNIST dataset. An important detail is that the input to the GRU at time $t$ is not the image $x_{t}$ , but rather the latent $\mathbf{Ub}_{t}$, encoded by the pre-trained backbone. After processing a sequence of digits ($x_{t-2},x_{t-1}$), the output state $o_{t}$ of the GRU was then used as a top-down context signal, in place of the one-hot context from the oracle case. We used a hidden state of size 128 for the GRU.
@@ 4.3.1 Loss function
The top-down network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function jointly across the following 5 conditions: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra). We use mean squared error as the loss criterion between the predicted (computed as in equation 2) and the target latent representations. Each mini-batch is sampled from dataset $D$, and training data for each condition are equally balanced in the mini-batch.
@@
[…]